OpenStudy (anonymous):

Help here please: Simplify the following expressions: a. cos(180° + x) + cos(180° - x) b. cos(126°)cos(36°) + sin(126°)sin(36°)

OpenStudy (anonymous):

Can you help me through a? I will solve b myself .

OpenStudy (experimentx):

you now cos C + cos D formula??

OpenStudy (campbell_st):

(a) cos(180)cos(x) - sin(180)sin(x) + cos(180)cos(x) + sin(180)sin(x) cos(180) = -1 so it simplifies to -cos(x) - cos(x) = -2cos(x)

OpenStudy (anonymous):

cos c =cos d?

OpenStudy (experimentx):

http://answers.yahoo.com/question/index?qid=20090209193628AAUvpWz you can use this formula formula or ... use above formula given by campbell_st

OpenStudy (campbell_st):

(b) will give cos(126 - 36) = cos(90) = 0

OpenStudy (experimentx):

above **method

OpenStudy (callisto):

cos(180° + x) = cox x in quadrant III =? cos(180° - x) = cox x in quadrant II =? Can you try these 2 first?

OpenStudy (anonymous):

so the first one is in Q3 and the second one in Q2, is it right?

OpenStudy (callisto):

yes, but what would it be? Sorry, typo : cox x -> cos x :S

OpenStudy (callisto):

For example cos(180° + x) = cos x in quadrant III = -cos x

OpenStudy (anonymous):

yeah, so both of them are -cos x right?

OpenStudy (callisto):

Yes~ so, what would you get for cos(180° + x) + cos(180° - x) ?

OpenStudy (anonymous):

-2cos x

OpenStudy (callisto):

yup~ first question is done :)

OpenStudy (anonymous):

I have a question, the formula that campbell used is the sum difference formula right?

OpenStudy (callisto):

it's compound angle formula

OpenStudy (anonymous):

that is weird, I didnt learn that one, I only learned the 6 sum difference formulas, and the double angle formula

OpenStudy (campbell_st):

question b is the expanded difference for cos... I recognised it straight away rewrote the question as the difference and got the solution.... its a working backwards question

OpenStudy (anonymous):

ok, let me see if I can manage question b

OpenStudy (anonymous):

is questino be a first pythagorian identity question?

OpenStudy (callisto):

Try to match this: cos (A-B) = cosAcosB + sinAsinB

OpenStudy (anonymous):

pythagorean*

OpenStudy (anonymous):

ok, one sec, I think i figured this out

OpenStudy (anonymous):